OpenStudy (anonymous):

find the perimeter of the right triangle given the length of one leg and the area a=6m area= 24m squared

OpenStudy (dumbcow):

area = 1/2*base*height base and height are the legs of triangle 24 = 1/2*6*b 24 = 3b 8 = b Then use pythagorean theorem to get c (hypotenuse) c^2 = 6^2 +8^2 c^2 = 100 c = 10 perimeter = 6+8+10